An air bubble of radius 3.5 cm is released from the mouth of a diver. At the diver's depth the pressure is 1.30 atm . (We'll learn in chapter 13 why the diver's depth affects the pressure.) As it rises to the surface, the air inside the bubble does 3.534 J of work. The bubble rises so quickly to the surface that this process can be treated as being adiabatic. If the bubble's temperature is initially the same as the diver's body (37∘C), what is its temperature when it reaches the surface? Answer is not 18.5 degrees celcius
Answer #1

chomy B Giren, initial radius of Sphere, Ri= 3.5x10-2 mm inition press are, Pa = 1.3x205-pa. initial temp. (Te) = 37° C = 310
Answer #3

soll? Initially, A Pressure (fi) = 1:30 atm = 1130 X o pa Temp (Ti)=370 = 310k. finally, Pressure (Pt) = 1 Atm = 1Xropa Temp
Answer #2

22 solution volume at the bubble at Divers depth . = V = 4103, 4x3114x (3.5X10 = 179.5x10-6 m3 so from pr=nRT → DR = PV = 11